OpenStudy (anonymous):

y=x2-14x+44 find the domain and range, and the vertex of the following parabola. Show all work? Help?

OpenStudy (anonymous):

@laughoutloud

OpenStudy (laughoutloud):

hey do want to know how to do it or just the answer either way doesn't matter to me

OpenStudy (anonymous):

Well, everytime I try to solve the vertex my teacher keeps saying its wrong. So I need help with that. Also with the whole domain and range thing.

OpenStudy (igreen):

@laughoutloud By the looks of his username, I think he wants to learn how to do it instead of just getting an answer.

OpenStudy (anonymous):

Yeah, I'd love to learn how to do it XD

OpenStudy (laughoutloud):

ok, to find the vertex you need to use the formula \[\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

Alright. So do I plug the x^2 into the 2a or the -b?

OpenStudy (laughoutloud):

that is going to help you find the x value

OpenStudy (anonymous):

and the -14 into the -b?

OpenStudy (laughoutloud):

you plug in \[\frac{ 14 }{ 2 }\]

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

That'd make X a 7?

OpenStudy (laughoutloud):

yes so x = 7 now you have to find the y value.

OpenStudy (laughoutloud):

What do you think you have to do?

OpenStudy (anonymous):

Alright, so I just plug 7 into all the X values

OpenStudy (laughoutloud):

yes

OpenStudy (anonymous):

Okay... Lemme solve it real quick

OpenStudy (anonymous):

y=(7)^2 - 14(7) + 44 y = 49 - 98 + 44 y = -49 + 44 y = -5 Is this correct?

OpenStudy (laughoutloud):

yes, I got the same

OpenStudy (laughoutloud):

so the vertex is?

OpenStudy (anonymous):

So, for the vertex it's (7, -5) right?

OpenStudy (laughoutloud):

yes

OpenStudy (anonymous):

Alright, how do we find the domain / range?

OpenStudy (laughoutloud):

the domain is going to be all of the x-values when this is graphed do you reach all of the x's eventually its kinda hard to explain

OpenStudy (laughoutloud):

That probably doesn't make sense

OpenStudy (anonymous):

Yeah... That's where I'm confused.

OpenStudy (laughoutloud):

Ok domain is all real

OpenStudy (anonymous):

Okay, sorry to trouble you more about the question... But sense I do all this stuff online and I'm just reading the stuff... How might I find the real numbers to explain to a teacher?

OpenStudy (laughoutloud):

sorry I'm not very good at explaining @carolinar7

OpenStudy (anonymous):

Alright, so.... Can you tell me how you got to your answer perhaps? @laughoutloud?